The Makers Co. has two producing departments: assembly and finishing. The company has been using a...

  1. The Makers Co. has two producing departments: assembly and finishing. The company has been using a plantwide predetermined overhead rate based on direct labor hours. The following estimates were made for the current year:

Assembly

Finishing

Total

Manufacturing Overhead

$200,000

$100,000

$300,000

Direct labor hours

40,000

35,000

75,000

Machine hours

5,000

16,000

21,000

Makers started and completed Job 1512 during the year. The job-order cost sheet indicated the following:

Materials Requisitioned

$18,000

Direct labor cost

16,000

Direct labor hours:

     Assembly

1,700 Hours

     Finishing

1,300 Hours

Machine Hours

     Assembly

1,000 Hours

     Finishing

700 Hours

A total of 2,000 units were produced on Job 1512.

Required:

1.

Assume that Makers uses a plant wide predetermined overhead based on direct labor hours. Calculate the total cost and the unit cost for each of the 2,000 units produced by Job 1512.

2.

Assume that Makers uses separate departmental overhead rates based upon direct labor hours for assembly and machine hours for finishing. Calculate the total cost and the unit cost for each of the 2,000 units produced by Job 1512.
Answer #1
Part-A
Plantwide Overhead Rate= Estimated Overhead/ Total Direct Labour Hour
=$300000/75000=$4 per Hour
Computaiton of Total Cost & Unit cost
Particular Amount
Material Cost $18,000.00
Direct Labour Cost $16,000.00
Manufacturing Overhead
Assembly ( 1700X$4) $6,800.00
Finishing ( 1300X$4) $5,200.00
Total Cost $46,000.00
Total Unit in Job 512 2000
Cost per Unit (46000/2000) $23.00
Part-b
Overhead Rate for Assembly Department = Estimated Overhead/ Total Direct Labour Hour
=200000/40000=$5per hour
. Overhead Rate for Finishing Department = Estimated Overhead/ Total Machine Hour
=100000/16000=$6.25
Computaiton of Total Cost & Unit cost
Particular Amount
Material Cost $18,000.00
Direct Labour Cost $16,000.00
Manufacturing Overhead
Assembly ( 1700X$5) $8,500.00
Finishing ( 700X$6.25) $4,375.00
Total Cost $46,875.00
Total Unit in Job 512 2000
Cost per Unit (46875/2000) $23.44
Let me know if any doubt Kindly mark with positive rating it would help me lot. Thanks in advance.